# Geometric Transformations on Null Curves in the Anti-de Sitter $3$-Space

### v1

@@ 3. The $\mathcal{T}$-Transform
Two star-shaped curves $\eta,\widetilde{\eta}:J\subseteq\mathbb{R}\longrightarrow\dot{\mathbb{R}}^{2}$ parameterized by the central affine arc length are said to be $\mathcal{T}$-transforms of each other [^3] Clearly, from Definition 3.1 it follows that the $\mathcal{T}$-transform is reciprocal. That is, if $\eta$ is a $\mathcal{T}$-transform of $\widetilde{\eta}$, then $\widetilde{\eta}$ is also a $\mathcal{T}$-transform of $\eta$. Equivalently, $\eta$ and $\widetilde{\eta}$ are $\mathcal{T}$-transforms of each other. We will use the different terms indistinctively. if ${\rm det}(\eta,\widetilde{\eta})$ is a nonzero constant.
@@
The $\mathcal{T}$-transform of star-shaped curves is a geometric transformation, in the sense that two star-shaped curves are $\mathcal{T}$-transforms of each other if the area of the triangles with vertices $O=(0,0)$, $\eta(s)$, and $\widetilde{\eta}(s)$ is constant for every $s\in J$. In Figures 2 and 3 we show a couple of examples of star-shaped curves and their corresponding $\mathcal{T}$-transforms and illustrate this geometric property.
Refer to caption   Refer to caption
Figure 2. The star-shaped curve $\eta_{+}$ of Figure 1, in blue and dotted in yellow, and the $\mathcal{T}$-transform $\widetilde{\eta}_{+}$ of $\eta_{+}$ in plain blue. The green triangles, which have the same area, are the ones with vertices $O$ (in black), $\eta_{+}(s)$ (in yellow) and $\widetilde{\eta}_{+}(s)$ (in blue). For each figure the triangle is shown at different values of the parameter $s\in J$.
Refer to caption   Refer to caption
Figure 3. The star-shaped curve $\eta_{-}$ of Figure 1, in red and dotted in yellow, and the $\mathcal{T}$-transform $\widetilde{\eta}_{-}$ of $\eta_{-}$ in plain red. The green triangles, which have the same area, are the ones with vertices $O$ (in black), $\eta_{-}(s)$ (in yellow) and $\widetilde{\eta}_{-}(s)$ (in blue). For each figure the triangle is shown at different values of the parameter $s\in J$.
@@
Let $\gamma,\widetilde{\gamma}:J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ be two null curves with associated pairs of cousins $(\eta_{+},\eta_{-})$ and $(\widetilde{\eta}_{+},\widetilde{\eta}_{-})$, respectively. We say that $\gamma$ and $\widetilde{\gamma}$ are $\mathcal{T}$-transforms of each other if $\eta_{+}$ is a $\mathcal{T}$-transform of $\widetilde{\eta}_{+}$ and $\eta_{-}$ is a $\mathcal{T}$-transform of $\widetilde{\eta}_{-}$.
@@
In Figures 2 and 3, we have shown the star-shaped $\mathcal{T}$-transforms $\widetilde{\eta}_{+}$ and $\widetilde{\eta}_{-}$ of the pair of cousins $(\eta_{+},\eta_{-})$ associated with the null curve $\gamma$ with constant bending $\kappa_{7,3}$ (see Figure 1). Hence, according to Definition 3.3, the null curve $\widetilde{\gamma}$ with associated pair of cousins $(\widetilde{\eta}_{+},\widetilde{\eta}_{-})$ is a $\mathcal{T}$-transform of $\gamma$. The curve $\widetilde{\gamma}$ will be shown in Figure 4.
Using the definition of the $\mathcal{T}$-transform on star-shaped curves we next find an equation satisfied by null curves that are $\mathcal{T}$-transform of each other.
@@
Let $\gamma,\widetilde{\gamma}:J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ be two null curves with associated pairs of cousins $(\eta_{+},\eta_{-})$ and $(\widetilde{\eta}_{+},\widetilde{\eta}_{-})$, respectively. If $\gamma$ and $\widetilde{\gamma}$ are $\mathcal{T}$-transforms of each other, then
(3.1) ${\rm det}(\eta_{+},\widetilde{\eta}_{+}){\rm det}(\eta_{-},\widetilde{\eta}_{-}^{\prime})-{\rm det}(\eta_{-},\widetilde{\eta}_{-}){\rm det}(\eta_{+},\widetilde{\eta}_{+}^{\prime})=\chi\,,$
@@
Let $\gamma$ and $\widetilde{\gamma}$ be two null curves that are $\mathcal{T}$-transforms of each other. It then follows from Definition 3.3 that $\eta_{+}$ is a $\mathcal{T}$-transform of $\widetilde{\eta}_{+}$ and $\eta_{-}$ is a $\mathcal{T}$-transform of $\widetilde{\eta}_{-}$. Moreover, Definition 3.1 then guarantees the existence of two nonzero constants $c_{\pm}\in\mathbb{R}$ such that ${\rm det}(\eta_{+},\widetilde{\eta}_{+})=c_{+}$ and ${\rm det}(\eta_{-},\widetilde{\eta}_{-})=c_{-}$, respectively. Then there exist smooth functions $f_{\pm}:J\subseteq\mathbb{R}\longrightarrow\mathbb{R}$ such that, respectively,
(3.2) $\widetilde{\eta}_{+}$ $=$ $\frac{1}{c_{+}}\left(\eta_{+}^{\prime}-f_{+}\eta_{+}\right),$
(3.3) $\widetilde{\eta}_{-}$ $=$ $\frac{1}{c_{-}}\left(\eta_{-}^{\prime}-f_{-}\eta_{-}\right).$
@@
(3.4) $\widetilde{\eta}_{+}^{\prime}=\frac{1}{c_{+}}\left(\left[\kappa+1-f_{+}^{\prime}\right]\eta_{+}-f_{+}\eta_{+}^{\prime}\right),$
where $\kappa$ is the bending of $\gamma$. Since $\eta_{+}$ and $\widetilde{\eta}_{+}$ are two star-shaped curves parameterized by the central affine arc length, ${\rm det}(\eta_{+},\eta_{+}^{\prime})=1$ and ${\rm det}(\widetilde{\eta}_{+},\widetilde{\eta}_{+}^{\prime})=1$. Thus, it follows from (3.2) and (3.4) that
11 11 $=$ ${\rm det}(\widetilde{\eta}_{+},\widetilde{\eta}_{+}^{\prime})=\frac{1}{c_{+}^{2}}{\rm det}(\eta_{+}^{\prime}-f_{+}\eta_{+},[\kappa+1-f_{+}^{\prime}]\eta_{+}-f_{+}\eta_{+}^{\prime})$
$=$ $\frac{1}{c_{+}^{2}}\left(f_{+}^{\prime}-\kappa-1+f_{+}^{2}\right).$
@@
(3.5) $f_{+}^{\prime}+f_{+}^{2}=\kappa+1+c_{+}^{2}\,.$
Combining (3.2), (3.4) and (3.5) we obtain that the canonical central affine frame field $\widetilde{F}_{+}$ along $\widetilde{\eta}_{+}$ is given in terms of the canonical central affine frame field $F_{+}$ along $\eta_{+}$ by
(3.6) $\widetilde{F}_{+}=\frac{1}{c_{+}}F_{+}\begin{pmatrix}-f_{+}&f_{+}^{2}-c_{+}^{2}\\ 1&-f_{+}\end{pmatrix}.$
@@
(3.7) $\widetilde{F}_{-}=\frac{1}{c_{-}}F_{-}\begin{pmatrix}-f_{-}&f_{-}^{2}-c_{-}^{2}\\ 1&-f_{-}\end{pmatrix}.$
@@
$\widetilde{F}_{+}^{-1}d\widetilde{F}_{+}$ $=$ $\frac{1}{c_{+}^{2}}\begin{pmatrix}-f_{+}&c_{+}^{2}-f_{+}^{2}\\ -1&-f_{+}\end{pmatrix}\left[F_{+}^{-1}dF_{+}\begin{pmatrix}-f_{+}&f_{+}^{2}-c_{+}^{2}\\ 1&-f_{+}\end{pmatrix}+\begin{pmatrix}-f_{+}^{\prime}&(f_{+}^{2})^{\prime}\\ 0&-f_{+}^{\prime}\end{pmatrix}\right]$
@@
(3.10) $\widetilde{\kappa}=2f_{+}^{2}-\kappa-2-2c_{+}^{2}\,.$
@@
(3.11) $\widetilde{\kappa}=2f_{-}^{2}-\kappa+2-2c_{-}^{2}\,.$
@@
$f_{+}=\pm\sqrt{2+c_{+}^{2}-c_{-}^{2}}\,\cosh(\phi)\,,\quad\quad\quad f_{-}=\sqrt{2+c_{+}^{2}-c_{-}^{2}}\,\sinh(\phi)\,.$
@@
±$\pm\sqrt{2+c_{+}^{2}-c_{-}^{2}}\,\sinh(\phi)\,\phi^{\prime}+(2+c_{+}^{2}-c_{-}^{2})\cosh^{2}(\phi)$ $=$ $\kappa+1+c_{+}^{2}\,,$
$\sqrt{2+c_{+}^{2}-c_{-}^{2}}\,\cosh(\phi)\,\phi^{\prime}+(2+c_{+}^{2}-c_{-}^{2})\sinh^{2}(\phi)$ $=$ $\kappa-1+c_{-}^{2}\,.$
Subtracting both equations we then see that $\phi^{\prime}=0$ and, hence, $\phi$ is constant. This implies that both $f_{+}$ and $f_{-}$ are also constant functions and, in addition, $f_{+}\neq f_{-}$.
@@
$\chi$ $=$ ${\rm det}(\eta_{+},\widetilde{\eta}_{+}){\rm det}(\eta_{-},\widetilde{\eta}_{-}^{\prime})-{\rm det}(\eta_{-},\widetilde{\eta}_{-}){\rm det}(\eta_{+},\widetilde{\eta}_{+}^{\prime})$
$=$ $\frac{1}{c_{+}}\left(\frac{-f_{-}}{c_{-}}\right)-\frac{1}{c_{-}}\left(\frac{-f_{+}}{c_{+}}\right)=\frac{f_{+}-f_{-}}{c_{+}c_{-}}\,,$
@@
$\chi={\rm det}(\eta_{+},\widetilde{\eta}_{+}){\rm det}(\eta_{-},\widetilde{\eta}_{-}^{\prime})-{\rm det}(\eta_{-},\widetilde{\eta}_{-}){\rm det}(\eta_{+},\widetilde{\eta}_{+}^{\prime})=\frac{f_{+}-f_{-}}{c_{+}c_{-}}=\frac{\ell}{c_{+}c_{-}}\,,$
@@ 3.1. Case $\chi\neq 0$
Let $\gamma:J=\mathbb{R}\longrightarrow{\rm AdS}$ be a null curve with bending $\kappa$ and $(F_{+},F_{-})$ be the spinor frame field along $\gamma$. If $\gamma$ has $\mathcal{T}$-transforms $\widetilde{\gamma}$ satisfying (3.1) with $\chi\neq 0$, then the bending $\kappa$ of $\gamma$ is constant and the $\mathcal{T}$-transforms $\widetilde{\gamma}$ are given by
@@
where $c_{+},c_{-}\in\mathbb{R}$ and $f_{+}\neq f_{-}$ are two constants such that
$f_{+}^{2}=\kappa+1+c_{+}^{2}\,,\quad\quad\quad f_{-}^{2}=\kappa-1+c_{-}^{2}\,.$
Moreover, the $\mathcal{T}$-transforms $\widetilde{\gamma}$ of $\gamma$ have constant bending $\widetilde{\kappa}=\kappa$.
@@
From the proof of Proposition 3.5, it follows that if $\chi\neq 0$, then both $f_{\pm}$ are constant functions. Hence, from the Riccati equation (3.5) (and the analogue one for $f_{-}$) we have
$f_{+}^{2}=\kappa+1+c_{+}^{2}\,,\quad\quad\quad f_{-}^{2}=\kappa-1+c_{-}^{2}\,.$
This proves that the bending $\kappa$ of $\gamma$ must be constant. Moreover, using this in the relation between $\kappa$ and the bending $\widetilde{\kappa}$ of $\widetilde{\gamma}$ given in (3.10), we conclude that $\widetilde{\kappa}=\kappa$.
Furthermore, since according to Theorem 2.3
$\widetilde{\gamma}=\widetilde{F}_{+}\widetilde{F}_{-}^{-1}\,,$
from the expression of the central affine frame field $\widetilde{F}_{+}$ given in (3.6) and the analogue one for $\widetilde{F}_{-}$ given in (3.7) it is a straightforward computation to obtain the explicit expression of the $\mathcal{T}$-transforms satisfying (3.1) with $\chi\neq 0$. ∎

### v2

@@ 3 The $\boldsymbol{\mathcal{T}}$-transform
Two star-shaped curves $\eta,\widetilde{\eta}\colon J\subseteq\mathbb{R}\longrightarrow\dot{\mathbb{R}}^{2}$ parameterized by the central affine arc length are said to be $\mathcal{T}$-transforms of each other [^3] Clearly, from Definition 3.1 it follows that the $\mathcal{T}$-transform is reciprocal. That is, if $\eta$ is a $\mathcal{T}$-transform of $\widetilde{\eta}$, then $\widetilde{\eta}$ is also a $\mathcal{T}$-transform of $\eta$. Equivalently, $\eta$ and $\widetilde{\eta}$ are $\mathcal{T}$-transforms of each other. We will use the different terms indistinctively. if $\det(\eta,\widetilde{\eta})$ is a nonzero constant.
@@
The $\mathcal{T}$-transform of star-shaped curves is a geometric transformation, in the sense that two star-shaped curves are $\mathcal{T}$-transforms of each other if the area of the triangles with vertices $O=(0,0)$, $\eta(s)$, and $\widetilde{\eta}(s)$ is constant for every $s\in J$. In Figures 2 and 3, we show a couple of examples of star-shaped curves and their corresponding $\mathcal{T}$-transforms and illustrate this geometric property.
Figure 2: The star-shaped curve $\eta_{+}$ of Figure 1, in blue and dotted in yellow, and the $\mathcal{T}$-transform $\widetilde{\eta}_{+}$ of $\eta_{+}$ in plain blue. The green triangles, which have the same area, are the ones with vertices $O$ (in black), $\eta_{+}(s)$ (in yellow) and $\widetilde{\eta}_{+}(s)$ (in blue). For each figure the triangle is shown at different values of the parameter $s\in J$.
Figure 3: The star-shaped curve $\eta_{-}$ of Figure 1, in red and dotted in yellow, and the $\mathcal{T}$-transform $\widetilde{\eta}_{-}$ of $\eta_{-}$ in plain red. The green triangles, which have the same area, are the ones with vertices $O$ (in black), $\eta_{-}(s)$ (in yellow) and $\widetilde{\eta}_{-}(s)$ (in red). For each figure the triangle is shown at different values of the parameter $s\in J$.
@@
Let $\gamma,\widetilde{\gamma}\colon J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ be two null curves with associated pairs of cousins $(\eta_{+},\eta_{-})$ and $(\widetilde{\eta}_{+},\widetilde{\eta}_{-})$, respectively. We say that $\gamma$ and $\widetilde{\gamma}$ are $\mathcal{T}$-transforms of each other if $\eta_{+}$ is a $\mathcal{T}$-transform of $\widetilde{\eta}_{+}$ and $\eta_{-}$ is a $\mathcal{T}$-transform of $\widetilde{\eta}_{-}$.
Figure 4: Left: The $\mathcal{T}$-transform (for $\chi=0$) $\widetilde{\gamma}=\mathcal{T}_{\xi,f}(\gamma)$ of the null curve $\gamma$ with constant bending $\kappa_{7,3}$ (see Figure 1). The parameter $\xi\neq 0$ of the $\mathcal{T}$-transform is $\xi=1.01$ and the transforming function $f$ is the solution of (3.11) with initial condition $0.1$. Right: The associated pair of star-shaped cousins $(\widetilde{\eta}_{+},\widetilde{\eta}_{-})$ in plain blue and red, respectively. The curves dotted in yellow are the corresponding $\mathcal{T}$-transforms, namely, $\eta_{+}$ and $\eta_{-}$ (cf. Figures 2 and 3).
@@
In Figures 2 and 3, we have shown the star-shaped $\mathcal{T}$-transforms $\widetilde{\eta}_{+}$ and $\widetilde{\eta}_{-}$ of the pair of cousins $(\eta_{+},\eta_{-})$ associated with the null curve $\gamma$ with constant bending $\kappa_{7,3}$ (see Figure 1). Hence, according to Definition 3.3, the null curve $\widetilde{\gamma}$ with associated pair of cousins $(\widetilde{\eta}_{+},\widetilde{\eta}_{-})$ is a $\mathcal{T}$-transform of $\gamma$. The curve $\widetilde{\gamma}$ is shown in Figure 4.
Using the definition of the $\mathcal{T}$-transform on star-shaped curves, we next find an equation satisfied by null curves that are $\mathcal{T}$-transform of each other.
@@
Let $\gamma,\widetilde{\gamma}\colon J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ be two null curves with associated pairs of cousins $(\eta_{+},\eta_{-})$ and $(\widetilde{\eta}_{+},\widetilde{\eta}_{-})$, respectively. If $\gamma$ and $\widetilde{\gamma}$ are $\mathcal{T}$-transforms of each other, then
$\det\bigl{(}\eta_{+},\widetilde{\eta}_{+}\bigr{)}\det\bigl{(}\eta_{-},\widetilde{\eta}_{-}^{\prime}\bigr{)}-\det\bigl{(}\eta_{-},\widetilde{\eta}_{-}\bigr{)}\det\bigl{(}\eta_{+},\widetilde{\eta}_{+}^{\prime}\bigr{)}=\chi$ (3.1)
@@
Let $\gamma$ and $\widetilde{\gamma}$ be two null curves that are $\mathcal{T}$-transforms of each other. It then follows from Definition 3.3 that $\eta_{+}$ is a $\mathcal{T}$-transform of $\widetilde{\eta}_{+}$ and $\eta_{-}$ is a $\mathcal{T}$-transform of $\widetilde{\eta}_{-}$. Moreover, Definition 3.1 then guarantees the existence of two nonzero constants $c_{\pm}\in\mathbb{R}$ such that ${\det(\eta_{+},\widetilde{\eta}_{+})=c_{+}}$ and $\det(\eta_{-},\widetilde{\eta}_{-})=c_{-}$, respectively.
Then there exist smooth functions $f_{\pm}\colon J\subseteq\mathbb{R}\longrightarrow\mathbb{R}$ such that, respectively,
$\widetilde{\eta}_{+}=\frac{1}{c_{+}}\bigl{(}\eta_{+}^{\prime}-f_{+}\eta_{+}\bigr{)},$ (3.2)
$\widetilde{\eta}_{-}=\frac{1}{c_{-}}\bigl{(}\eta_{-}^{\prime}-f_{-}\eta_{-}\bigr{)}.$ (3.3)
@@
$\widetilde{\eta}_{+}^{\prime}=\frac{1}{c_{+}}\bigl{(}\bigl{[}\kappa+1-f_{+}^{\prime}\bigr{]}\eta_{+}-f_{+}\eta_{+}^{\prime}\bigr{)},$ (3.4)
where $\kappa$ is the bending of $\gamma$. Since $\eta_{+}$ and $\widetilde{\eta}_{+}$ are two star-shaped curves parameterized by the central affine arc length, $\det\bigl{(}\eta_{+},\eta_{+}^{\prime}\bigr{)}=1$ and $\det\bigl{(}\widetilde{\eta}_{+},\widetilde{\eta}_{+}^{\prime}\bigr{)}=1$. Thus, it follows from (3.2) and (3.4) that
$1=\det\bigl{(}\widetilde{\eta}_{+},\widetilde{\eta}_{+}^{\prime}\bigr{)}=\frac{1}{c_{+}^{2}}\det\bigl{(}\eta_{+}^{\prime}-f_{+}\eta_{+},\bigl{[}\kappa+1-f_{+}^{\prime}\bigr{]}\eta_{+}-f_{+}\eta_{+}^{\prime}\bigr{)}=\frac{1}{c_{+}^{2}}\bigl{(}f_{+}^{\prime}-\kappa-1+f_{+}^{2}\bigr{)}.$
@@
$f_{+}^{\prime}+f_{+}^{2}=\kappa+1+c_{+}^{2}.$ (3.5)
Combining (3.2), (3.4) and (3.5), we obtain that the canonical central affine frame field $\widetilde{F}_{+}$ along $\widetilde{\eta}_{+}$ is given in terms of the canonical central affine frame field $F_{+}$ along $\eta_{+}$ by
$\widetilde{F}_{+}=\frac{1}{c_{+}}F_{+}\begin{pmatrix}-f_{+}&f_{+}^{2}-c_{+}^{2}\\ 1&-f_{+}\end{pmatrix}.$ (3.6)
@@
$\widetilde{F}_{-}=\frac{1}{c_{-}}F_{-}\begin{pmatrix}-f_{-}&f_{-}^{2}-c_{-}^{2}\\ 1&-f_{-}\end{pmatrix}.$ (3.7)
@@
$\widetilde{F}_{+}^{-1}{\rm d}\widetilde{F}_{+}$ ${}=\frac{1}{c_{+}^{2}}\begin{pmatrix}-f_{+}&c_{+}^{2}-f_{+}^{2}\\ -1&-f_{+}\end{pmatrix}\biggl{[}F_{+}^{-1}{\rm d}F_{+}\begin{pmatrix}-f_{+}&f_{+}^{2}-c_{+}^{2}\\ 1&-f_{+}\end{pmatrix}+\begin{pmatrix}-f_{+}^{\prime}&(f_{+}^{2})^{\prime}\\ 0&-f_{+}^{\prime}\end{pmatrix}\biggr{]}$
@@
$\widetilde{\kappa}=2f_{+}^{2}-\kappa-2-2c_{+}^{2}.$ (3.8)
@@
$\widetilde{\kappa}=2f_{-}^{2}-\kappa+2-2c_{-}^{2}.$ (3.9)
@@
$f_{+}=\pm\sqrt{2+c_{+}^{2}-c_{-}^{2}}\cosh(\phi),\qquad f_{-}=\sqrt{2+c_{+}^{2}-c_{-}^{2}}\sinh(\phi).$
@@
±$\begin{split}&\pm\sqrt{2+c_{+}^{2}-c_{-}^{2}}\sinh(\phi)\phi^{\prime}+\bigl{(}2+c_{+}^{2}-c_{-}^{2}\bigr{)}\cosh^{2}(\phi)=\kappa+1+c_{+}^{2},\\ &\sqrt{2+c_{+}^{2}-c_{-}^{2}}\cosh(\phi)\phi^{\prime}+\bigl{(}2+c_{+}^{2}-c_{-}^{2}\bigr{)}\sinh^{2}(\phi)=\kappa-1+c_{-}^{2}.\end{split}$
Subtracting both equations, we then see that $\phi^{\prime}=0$ and, hence, $\phi$ is constant. This implies that both $f_{+}$ and $f_{-}$ are also constant functions and, in addition, $f_{+}\neq f_{-}$.
@@
$\chi$ ${}=\det(\eta_{+},\widetilde{\eta}_{+})\det\bigl{(}\eta_{-},\widetilde{\eta}_{-}^{\prime}\bigr{)}-\det(\eta_{-},\widetilde{\eta}_{-})\det\bigl{(}\eta_{+},\widetilde{\eta}_{+}^{\prime}\bigr{)}$
${}=\frac{1}{c_{+}}\biggl{(}\frac{-f_{-}}{c_{-}}\biggr{)}-\frac{1}{c_{-}}\biggl{(}\frac{-f_{+}}{c_{+}}\biggr{)}=\frac{f_{+}-f_{-}}{c_{+}c_{-}},$
@@
$\chi=\det(\eta_{+},\widetilde{\eta}_{+})\det\bigl{(}\eta_{-},\widetilde{\eta}_{-}^{\prime}\bigr{)}-\det(\eta_{-},\widetilde{\eta}_{-})\det\bigl{(}\eta_{+},\widetilde{\eta}_{+}^{\prime}\bigr{)}=\frac{f_{+}-f_{-}}{c_{+}c_{-}}=\frac{\ell}{c_{+}c_{-}},$
@@ 3.1 Case $\boldsymbol{\chi\neq 0}$
Let $\gamma\colon J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ be a null curve with bending $\kappa$ and $(F_{+},F_{-})$ be the spinor frame field along $\gamma$. If $\gamma$ has $\mathcal{T}$-transforms $\widetilde{\gamma}$ satisfying (3.1) with $\chi\neq 0$, then the bending $\kappa$ of $\gamma$ is constant and the $\mathcal{T}$-transforms $\widetilde{\gamma}$ are given by
@@
where $c_{+},c_{-}\in\mathbb{R}$ and $f_{+}\neq f_{-}$ are two constants such that $f_{+}^{2}=\kappa+1+c_{+}^{2}$, $f_{-}^{2}=\kappa-1+c_{-}^{2}$. Moreover, the $\mathcal{T}$-transforms $\widetilde{\gamma}$ of $\gamma$ have constant bending $\widetilde{\kappa}=\kappa$.
@@
From the proof of Proposition 3.5, it follows that if $\chi\neq 0$, then both $f_{\pm}$ are constant functions. Hence, from the Riccati equation (3.5) (and the analogue one for $f_{-}$), we have $f_{+}^{2}=\kappa+1+c_{+}^{2}$, $f_{-}^{2}=\kappa-1+c_{-}^{2}$. This proves that the bending $\kappa$ of $\gamma$ must be constant. Moreover, using this in the relation between $\kappa$ and the bending $\widetilde{\kappa}$ of $\widetilde{\gamma}$ given in (3.8), we conclude that $\widetilde{\kappa}=\kappa$.
Furthermore, since according to Theorem 2.4 $\widetilde{\gamma}=\widetilde{F}_{+}\widetilde{F}_{-}^{-1}$, from the expression of the central affine frame field $\widetilde{F}_{+}$ given in (3.6) and the analogue one for $\widetilde{F}_{-}$ given in (3.7) it is a straightforward computation to obtain the explicit expression of the $\mathcal{T}$-transforms satisfying (3.1) with $\chi\neq 0$. ∎